OpenStudy (anonymous):

What is the direct variation equation if f(x) varies directly with x and f(x) = –9 when x = 3? f(x) = -1/3x f(x) = –3x f(x) = –6x f(x) = –12x

OpenStudy (flcon04):

Any ideas to what the answer is?

OpenStudy (anonymous):

I think it's D

OpenStudy (flcon04):

Its not D try swapping in the number into the equation.

OpenStudy (anonymous):

I got it it's C

OpenStudy (flcon04):

one second ill double check.

OpenStudy (flcon04):

wait I mean B.

OpenStudy (anonymous):

I knew it lol I had to think for a sec. Thanks

OpenStudy (flcon04):

I got confused halfway thru lol.

OpenStudy (trojanpoem):

\[f(x) \alpha x \] f(x) = kx (k const) -9 = 3k , k = -3 f(x) = -3x